OpenStudy (mathmath333):

remainder for 4^875/17

ganeshie8 (ganeshie8):

any ideas ?

OpenStudy (ikram002p):

4^17= 4 mod 17 xD ?!

ganeshie8 (ganeshie8):

haha Fermat is not efficient here as it leaves a very huge remainder

OpenStudy (ikram002p):

so 4^18=-1 mod 17

OpenStudy (ikram002p):

:P no its not :P :P

ganeshie8 (ganeshie8):

@mathmath333 u have any ideas ?

OpenStudy (mathmath333):

2^1750

OpenStudy (ikram002p):

ok lol something else xD 4 = 4 mod 17 4^2= 16 mod 17 4^2 =-1 mod 17 ?! is this more help ?

OpenStudy (mathmath333):

book gives the ans as 13

ganeshie8 (ganeshie8):

you're on right track @mathmath333

OpenStudy (mathmath333):

2^(3*586+2) ?

ganeshie8 (ganeshie8):

try this : \[ \large 4^{875} = 4^{2\times 437 + 1} = 4*16^{437} \equiv 4*(-1)^{437} \equiv -4 \equiv -4+17\equiv 13\]

OpenStudy (08surya):